If x, y and z are consecutive terms in a geometric sequence, then y over z is equal to z over y. Show that the following numbers are consecutively in a geometric sequence:
A) 5, square root of 35 and 7
B) a, square root of ab and b

Answer by stanbon(75887)

If x, y and z are consecutive terms in a geometric sequence, then y over z is equal to z over y. Show that the following numbers are consecutively in a geometric sequence:
A) 5, square root of 35 and 7
r = sqrt(35)/5
r = 7/sqrt(35) = 7*sqrt(35)/35 = sqrt(35)/5
Since the ratios are the same, the sequence is geometric.
-----------------------------------
B) a, square root of ab and b
r = sqrt(ab)/a
r = b/sqrt(ab) = b*sqrt(ab)/ab = sqrt(ab)/a
Since the ratios are the same, the sequence is geometric.
-----
Cheers,
Stan H.
------------